### v2

@@ Abstract
Hardware trojan detection requires both accurate identification and interpretable explanations that enable hardware security engineers to understand and act on detection results. While machine learning approaches show promise, existing explainability methods developed for general domains like image classification may not provide the actionable insights that hardware engineers need. A question remains: How do domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution techniques compare for hardware security applications?
This work compares three categories of explainability for gate-level hardware trojan detection on the Trust-Hub benchmark dataset: (1) domain-aware property-based analysis of 31 circuit-specific features derived from gate fanin patterns, flip-flop distances, and primary I/O connectivity; (2) model-agnostic case-based reasoning using k-nearest neighbors for precedent-based explanations; and (3) model-agnostic feature attribution methods (Local Interpretable Model-agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), gradient) that provide generic importance scores without circuit-level context.
The findings show that different explainability approaches offer distinct advantages for hardware security practitioners. The domain-aware property-based method analyzes 31 circuit properties, enabling engineers to understand detection decisions through familiar concepts like "high fanin complexity near outputs indicates potential trojan triggers." Case-based reasoning (k-nearest neighbors) achieves 97.4% correspondence between predictions and training exemplars, providing model-agnostic justifications grounded in precedent that require minimal domain interpretation. In contrast, while LIME and SHAP provide mathematically rigorous feature attributions with strong inter-method correlation ($\rho$=0.94, p<<0.001), they produce generic feature importance scores that lack circuit-level context for validation or remediation planning. Detection performance using eXtreme Gradient Boosting (XGBoost) classification with optimized threshold achieves 46.15% precision and 52.17% recall on 11,392 test samples, representing a 9-fold improvement over prior work (Hasegawa et al.: 5.13% precision) while reducing false positive rates from approximately 5.6% to 0.25%.
Additionally, gradient-based feature attribution (Simonyan et al., 2013), while 481 times faster than SHAP, provides similar domain-opaque insights as other model-agnostic methods, confirming that computational efficiency alone cannot substitute for domain-aware interpretability.
@@ 1 Introduction
When a machine learning system flags a gate in a billion-transistor integrated circuit as a potential hardware trojan, what should the security engineer do next? Generic explanations like "feature importance: 0.73" provide little guidance for validating the detection or planning remediation. This explainability gap represents a barrier to deploying ML-based trojan detection in practice: without actionable interpretations grounded in circuit design principles, engineers cannot distinguish meaningful alerts from false positives, cannot verify suspicious structural patterns against domain knowledge, and cannot translate detection results into concrete security actions.
@@
Hardware trojans are malicious circuit modifications introduced during design or fabrication that pose unique challenges compared to software vulnerabilities. Unlike software patches, hardware trojans are permanent components of silicon. Unlike software bugs triggered by common code paths, hardware trojans activate through rare input combinations designed to evade conventional testing (as few as 1 in $2^{64}$ possible states). Unlike software exploits visible through runtime monitoring, hardware trojans operate at the gate-level where they can compromise system confidentiality, integrity, or availability with no software-visible trace. As infrastructure, financial systems, and defense applications depend on integrated circuits fabricated through complex global supply chains involving potentially untrusted facilities, hardware trojan detection capabilities have become important for supply chain security and system integrity.
@@
Machine learning approaches have demonstrated promise for hardware trojan detection by identifying suspicious circuit structures through statistical analysis of gate-level netlists. However, existing ML-based detection methods face a challenge: opaque-box models provide predictions without interpretable justifications that hardware security engineers can understand and act upon. While XAI techniques have been developed to interpret ML predictions, these methods were primarily designed for general domains like image classification and natural language processing. This raises the question: Do model-agnostic explainability techniques developed for general ML applications provide the actionable insights that hardware security practitioners need, or do domain-aware methods tailored to circuit characteristics offer interpretability?
@@
Explainability is important in hardware trojan detection because security engineers need actionable insights: not just feature importance scores, but interpretations grounded in circuit design principles that enable validation and remediation. When a detection system flags a gate as suspicious, engineers need to understand the finding in terms they can verify: "This gate exhibits high fanin complexity (12 inputs at 2 levels) near a primary output, matching the signature of trojan trigger circuits." Generic explanations like "feature X has importance 0.73" provide limited actionability for domain experts.
@@
This distinction between domain-aware and model-agnostic explainability represents a tradeoff in XAI system design. Model-agnostic methods like LIME [1] and SHAP [2] offer mathematical rigor and applicability across any ML model, explaining predictions through feature perturbation and game-theoretic attribution. These methods have been applied effectively to image classification and text analysis where generic feature importance suffices. However, in specialized domains like hardware security, model-agnostic methods may lack the domain context necessary for practitioners to trust and act on explanations.
@@
Domain-aware explainability methods, in contrast, incorporate domain knowledge directly into the explanation framework. For hardware trojan detection, this means analyzing circuit-specific properties like gate fanin patterns, flip-flop connectivity, and proximity to primary I/O, characteristics that hardware engineers routinely use to reason about circuit behavior. While domain-aware methods require domain expertise to design, they align with existing practitioner knowledge and reasoning patterns. Yet prior work has not empirically evaluated whether this theoretical advantage translates to practical benefits for hardware security applications.
@@
    Domain-Aware Property Analysis: A property-based explainability method is developed and evaluated that analyzes 31 circuit-specific features grounded in hardware design principles. This domain-aware approach provides explanations like "high Logic Gate Fanin (LGFi)=12 at 2 levels upstream combined with low Flip-Flop Output distance (FFo)=1 matches rare-event trigger patterns," enabling hardware engineers to validate findings through familiar circuit analysis concepts.
@@
    Case-Based Reasoning with k-Nearest Neighbors: An inherently interpretable prediction method using k-nearest neighbors (k=5) achieves 97.4% correspondence between predictions and training exemplars. This approach provides justifications grounded in precedent: "This gate’s feature profile matches 4 out of 5 similar training cases that were trojans," enabling engineers to validate detections through comparison with known patterns.
@@
    Comparison with Model-Agnostic Baselines: Domain-aware and inherently interpretable methods are compared against model-agnostic techniques (LIME, SHAP) that provide generic feature importance without circuit-level context. While LIME and SHAP exhibit strong inter-method correlation ($\rho$=0.94, p<<0.001) validating their mathematical consistency, generic feature rankings provide limited actionability for hardware security practitioners. Gradient-based attribution (Simonyan et al., 2013) is also evaluated for computational efficiency.
@@
    Improved Detection Performance with XGBoost: The XGBoost-based detector achieves 46.15% precision (95% CI: [42.06%, 50.24%]) at 52.17% recall (95% CI: [45.65%, 58.70%]) on the Trust-Hub benchmark dataset, representing a 9-fold improvement in precision over the Hasegawa et al. baseline (5.13% precision, 82.6% recall). This improvement addresses a key limitation for practical deployment where engineers must investigate flagged gates.
@@ 2.1 Hardware Trojans
A hardware trojan circuit taxonomy is shown in Figure 1. A hardware trojan circuit has two primary parts: the trigger, which activates the trojan, and the payload, which delivers the exploit. Triggers may be classified as digital or analog. Off-die sensors may trigger analog triggers. Rare values, k-bit synchronous or asynchronous delays, and rare sequences may activate triggers. Payloads may also be digital or analog. Digital payloads could alter control, status, or data lines. Analog payloads could exploit the circuit by introducing leakage, bridging, delay, or unintended activity[3].
@@
Figure 2 shows a specific implementation of a confidential hardware trojan using logic gates within an Integrated Circuit (IC) design file netlist. The dashed lines represent additional levels of logic gates. The trigger is conditional on input patterns within the IC, which eventually originate from the primary inputs. The logical-and gate requires that all the inputs are logically true before triggering the payload. To be difficult to detect or accidentally trigger during IC test mode, the hardware trojan trigger should be a rare event condition. Given $n$ inputs to a combinational IP circuit, the rarest condition would be one of $2^{n}$ inputs. Once the hardware trojan is triggered, the multiplexor circuit within the payload switches from normal operation to leaking sensitive information on the primary output pin of the circuit where the attacker can observe it.
This rare event hardware trojan model allows for detection through static circuit analysis by examining each net or group of logic gates within an IP circuit primary for input or fanin complexity. The greater the fanin, the higher the probability that a subcircuit is a potential suspect for a hardware trojan. Additionally, payloads must propagate their sensitive information to the primary output. The rare event fanin and primary output payload concepts form the basis of trojan analysis of several methodologies [3, 4, 5, 6].
@@
Hasegawa et al. introduced a method of detecting hardware trojans from a gate-level netlist [4, 5]. Five features from each net are used to classify trojans. (a) Logic Gate Fanins: LGFi represent the number of inputs to the logic gates two levels upstream from a net. (b) Flip-Flop Input (FFi) represents the number of upstream logic levels to a flip-flop. (c) Flip-Flop Output: FFo represents the number of logic levels downstream to a flip-flop. (d) Primary Input (PI) which represents the number of logic levels upstream to the closest primary input. (e) Primary Output (PO) represents the number of logic levels downstream to the closest primary output. In the work, the authors also used three training strategies, due to the imbalance of the trojan to non-trojan training data ratio: no weighting, static weighting, and dynamic weighting. Dynamic weighting performed best. Other works address the imbalance in other means, such as synthetic method over-sampling techniques [7]. These optimized methods still do not provide a better understanding of the ML results.
@@ 2.2 Trust-Hub Benchmark Dataset
The Trust-Hub benchmark dataset [8, 9, 10, 6] provides gate-level netlists with known trojan insertions for various circuit types. This work uses 30 circuits comprising RS232 communication peripherals and International Symposium on Circuits and Systems (ISCAS) combinational benchmarks.
@@ 3.1 Machine Learning for Hardware Trojan Detection
Precision computed from their published confusion matrices is 5.13%, indicating that approximately 95% of gates flagged as trojans are false positives. With a dataset imbalance of approximately 275:1 (benign to trojan), this translates to a false positive rate of approximately 5.6%, meaning that while most benign gates are correctly classified, the sheer volume of false positives (about 20× the number of actual trojans) limits practical utility for real-world deployment where security engineers must manually inspect falsely flagged gates. This work addresses this limitation by achieving 46.15% precision (95% bootstrap CI: [7.14%, 13.53%] at baseline threshold 0.5, improving to 46.15% at optimized threshold 0.99), a 9-fold improvement that reduces the false positive burden.
@@ 3.1.2 Recent Advances
In contrast, this work performs gate-level (node-level) trojan localization on heavily imbalanced datasets (approximately 275:1 ratio of benign to trojan gates), directly identifying suspicious gates to enable targeted remediation. While HT-Pred achieves higher overall accuracy on the circuit-level task, the gate-level localization provides the specific information needed for practical trojan removal or mitigation.
@@ 3.1.4 Class Imbalance Handling in Hardware Trojan Detection
Hardware trojan detection faces extreme class imbalance: in typical gate-level netlists, benign gates outnumber trojan gates by ratios exceeding 100:1, often reaching 275:1 as observed in the Trust-Hub benchmark dataset. Standard machine learning classifiers trained on such imbalanced data exhibit bias toward the majority class, achieving high overall accuracy while failing to detect minority-class trojans.
@@
This work employs XGBoost’s scale_pos_weight parameter to implement cost-sensitive learning following King and Zeng’s weighting approach [12]. By setting the weight ratio equal to the class imbalance ratio ($\frac{N_{benign}}{N_{trojan}}$), the classifier assigns higher loss penalties to trojan misclassifications, encouraging the model to learn discriminative patterns despite extreme rarity. This approach achieves 52.17% recall on 275:1 imbalanced test data, demonstrating that proper imbalance handling is essential for practical trojan detection deployment.
@@ 3.2.2 XAI for Hardware Security
The application of explainable AI to hardware trojan detection remains largely unexplored in published literature. A systematic literature search across IEEE Xplore, ACM Digital Library, and Google Scholar using query terms: “explainable AI hardware trojan,” “interpretable machine learning hardware security,” “XAI hardware trojan detection,” and “explainability gate-level netlist analysis.” The search covered publications from 2016 to 2024 (post-LIME and SHAP introduction) and yielded zero peer-reviewed papers systematically evaluating XAI methods for hardware trojan detection.
Most prior work on ML-based trojan detection uses opaque-box models without explainability analysis [4, 5, 11, 7]. When explanations are provided, they typically consist of simple feature importance rankings extracted from tree-based models (e.g., Random Forest feature importances) without rigorous evaluation of explanation quality, consistency across methods, or actionability for security practitioners. No prior work has compared domain-aware versus model-agnostic explainability techniques, assessed explanation consistency through inter-method correlation, or quantified computational costs of explanation generation at the gate-level.
No prior work has systematically evaluated XAI methods for hardware trojan detection, despite the importance of explainability for security applications. Hardware security engineers must validate detection results through manual inspection and design remediation strategies based on identified vulnerabilities. Generic feature importance scores like “LGFi has weight 0.73” provide limited actionability compared to domain-grounded explanations like “high fanin complexity (12 inputs at 2 levels) near primary output matches rare-event trigger signature.”

### v3

@@ Abstract
Hardware trojans are malicious circuits which compromise the functionality and security of an integrated circuit (IC). These circuits are manufactured directly into the silicon and cannot be fixed by security patches like software. The solution would require a costly product recall by replacing the IC and hence, early detection in the design process is essential. Hardware detection at best provides statistically based solutions with many false positives and false negatives. These detection methods require more thorough explainable analysis to filter out false indicators.
Existing explainability methods developed for general domains like image classification may not provide the actionable insights that hardware engineers need. A question remains: How do domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution techniques compare for hardware security applications?
This work compares three categories of explainability for gate-level hardware trojan detection on the Trust-Hub benchmark dataset: (1) domain-aware property-based analysis of 31 circuit-specific features derived from gate fanin patterns, flip-flop distances, and primary Input/Output (I/O) connectivity; (2) model-agnostic case-based reasoning using k-nearest neighbors for precedent-based explanations; and (3) model-agnostic feature attribution methods (Local Interpretable Model-agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), gradient) that provide generic importance scores without circuit-level context.
The findings show that different explainability approaches offer distinct advantages for hardware security practitioners. The domain-aware property-based method analyzes 31 circuit properties, enabling engineers to understand detection decisions through familiar concepts like “high fanin complexity near outputs indicates potential trojan triggers.” Case-based reasoning (k-nearest neighbors) achieves 95.82% correspondence between predictions and training exemplars, providing model-agnostic justifications grounded in precedent that require minimal domain interpretation. In contrast, while LIME and SHAP provide mathematically rigorous feature attributions with strong inter-method Spearman rank correlation ($\rho_{s}=1.000$, $p<0.001$), they produce generic feature importance scores that lack circuit-level context for validation or remediation planning. Detection performance using eXtreme Gradient Boosting (XGBoost) classification with optimized threshold achieves 72.06% precision (95% CI: [61.11%, 82.54%]) and 62.03% recall (95% CI: [51.22%, 72.58%]) on 11,392 test samples, representing a 14-fold improvement at the optimized threshold over prior work (Hasegawa et al.: 5.13% precision; 2.3-fold at the default threshold of 0.5) while reducing false positive rates from approximately 50.5% to 0.17%.
Additionally, gradient-based feature attribution (Simonyan et al., 2013), with a speedup factor of 7 over SHAP, provides similar domain-opaque insights as other model-agnostic methods, confirming that computational efficiency alone cannot substitute for domain-aware interpretability.
@@ 1 Introduction
When a machine learning system flags a gate in a billion-transistor integrated circuit as a potential hardware trojan, what should the security engineer do next? Generic explanations like “feature importance: 0.73” provide little guidance for validating the detection or planning remediation. This explainability gap represents a barrier to deploying ML-based trojan detection in practice: without actionable interpretations grounded in circuit design principles, engineers cannot distinguish meaningful alerts from false positives, cannot verify suspicious structural patterns against domain knowledge, and cannot translate detection results into concrete security actions.
@@
Hardware trojans are malicious circuit modifications introduced during design or fabrication that pose unique challenges compared to software vulnerabilities. Unlike software patches, hardware trojans are permanent components of silicon. Unlike software bugs triggered by common code paths, hardware trojans activate through rare input combinations designed to evade conventional testing (as few as 1 in $2^{64}$ possible states). Unlike software exploits visible through runtime monitoring, hardware trojans operate at the gate-level where they can compromise system confidentiality, integrity, or availability with no software-visible trace.
@@
Machine learning approaches have demonstrated promise for hardware trojan detection by identifying suspicious circuit structures through statistical analysis of gate-level netlists. However, existing ML-based detection methods face a challenge: opaque-box models provide predictions without interpretable justifications that hardware security engineers can understand and act upon. While XAI techniques have been developed to interpret ML predictions, these methods were primarily designed for general domains like image classification and natural language processing. A model-agnostic XAI method is one that can be applied to any ML model without requiring knowledge of the model’s internal structure while a domain-aware method incorporates domain knowledge into the explanation process such as gate-level connectivity and circuit topology in hardware trojans. This raises the question: Do model-agnostic explainability techniques developed for general ML applications provide the actionable insights that hardware security practitioners need, or do domain-aware methods tailored to circuit characteristics offer interpretability?
@@
Explainability is important in hardware trojan detection because security engineers need actionable insights: not just feature importance scores, but interpretations grounded in circuit design principles that enable validation and remediation. When a detection system flags a gate as suspicious, engineers need to understand the finding in terms they can verify: “This gate exhibits high fanin complexity (12 inputs at 2 levels) near a primary output, matching the signature of trojan trigger circuits.” Generic explanations like “feature X has importance 0.73” provide limited actionability for domain experts.
@@
This distinction between domain-aware and model-agnostic explainability represents a tradeoff in XAI system design. Model-agnostic methods like LIME [1] and SHAP [2] offer mathematical techniques and applicability across any ML model, explaining predictions through feature perturbation and game-theoretic attribution. These methods have been applied effectively to image classification and text analysis where generic feature importance suffices. However, in specialized domains like hardware security, model-agnostic methods may lack the domain context necessary for practitioners to trust and act on explanations.
@@
Domain-aware explainability methods, in contrast, incorporate domain knowledge directly into the explanation framework. For hardware trojan detection, this means analyzing circuit-specific properties like logic gate fanin patterns, flip-flop connectivity, and proximity to primary I/O, characteristics that hardware engineers routinely use to reason about circuit behavior. While domain-aware methods require domain expertise to design, they align with existing practitioner knowledge and reasoning patterns. Yet prior work has not empirically evaluated whether this theoretical advantage translates to practical benefits for hardware security applications.
@@
    Domain-Aware Property Analysis: A property-based explainability method is developed and evaluated that analyzes 31 gate-level circuit-specific features grounded in hardware design principles. This domain-aware approach provides explanations like, “High Logic Gate Fanin (LGFi)=12 at 2 levels upstream combined with low Flip-Flop Output distance (FFo)=1 matches rare-event trigger patterns,” enabling hardware engineers to validate findings through familiar circuit analysis concepts.
@@
    Case-Based Reasoning with k-Nearest Neighbors: An inherently interpretable prediction method using k-nearest neighbors (k=5) achieves 95.82% correspondence between predictions and training exemplars. This approach provides justifications grounded in precedent:“This gate’s feature profile matches 4 out of 5 similar training cases that were trojans,” enabling engineers to validate detections through comparison with known patterns.
@@
    Comparison with Model-Agnostic Baselines: Domain-aware and inherently interpretable methods are compared against model-agnostic techniques (LIME, SHAP) that provide generic feature importance without circuit-level context. While LIME and SHAP exhibit strong inter-method Spearman rank correlation ($\rho_{s}=1.000$, $p<0.001$, where $\rho_{s}$ is the Spearman correlation coefficient and $p$ is the p-value) validating their mathematical consistency, generic feature rankings provide limited actionability for hardware security practitioners. Gradient-based attribution (Simonyan et al., 2013) is also evaluated for computational efficiency.
@@
    Improved Detection Performance with XGBoost: The XGBoost-based detector achieves 72.06% precision (95% CI: [61.11%, 82.54%]) at 62.03% recall (95% CI: [51.22%, 72.58%]) on the Trust-Hub benchmark dataset, representing a 14-fold improvement in precision at the optimized threshold over the Hasegawa et al. baseline (5.13% precision, 82.6% recall), or 2.3-fold at the default threshold. This improvement addresses a key limitation for practical deployment where engineers must investigate flagged gates.
@@ 2.1 Hardware Trojans
A hardware trojan circuit taxonomy is shown in Figure 1. A hardware trojan circuit has two primary parts: the trigger, which activates the trojan, and the payload, which delivers the exploit. Triggers may be classified as digital or analog. Off-die sensors may activate analog triggers. Rare values, k-bit synchronous or asynchronous delays, and rare sequences may activate triggers. Payloads may also be digital or analog. Digital payloads could alter control, status, or data lines. Analog payloads could exploit the circuit by introducing leakage, bridging, delay, or unintended activity[3].
@@
Figure 2 shows a specific implementation of a confidentiality hardware trojan using logic gates within an Integrated Circuit (IC) design file netlist. The dashed lines represent additional levels of logic gates. The trigger is conditional on input patterns within the IC, which eventually originate from the primary inputs. The logical-and gate requires that all the inputs are logically true before triggering the payload. To be difficult to detect or accidentally trigger during IC test mode, the hardware trojan trigger should be a rare event condition. Given $n$ inputs to a combinational IP circuit, the rarest condition would be one of $2^{n}$ inputs. Once the hardware trojan is triggered, the multiplexor circuit within the payload switches from normal operation to leaking sensitive information on the primary output pin of the circuit where the attacker can observe it.
This rare event hardware trojan model allows for detection through static circuit analysis by examining each net or group of logic gates within an IP circuit, primarily for input or fanin complexity. The greater the fanin, the higher the probability that a subcircuit is a potential suspect for a hardware trojan. Additionally, payloads must propagate their sensitive information to the primary output. The rare event fanin and primary output payload concepts form the basis of trojan analysis of several methodologies [3, 4, 5, 6].
@@
Hasegawa et al. introduced a method of detecting hardware trojans from a gate-level netlist [4, 5]. Five features from each net are used to classify trojans. (a) Logic Gate Fanins: LGFi represents the number of inputs to the logic gates two levels upstream from a net. (b) Flip-Flop Input (FFi) represents the number of upstream logic levels to a flip-flop. (c) Flip-Flop Output: FFo represents the number of logic levels downstream to a flip-flop. (d) Primary Input (PI), which represents the number of logic levels upstream to the closest primary input. (e) Primary Output (PO) represents the number of logic levels downstream to the closest primary output. In the work, the authors also used three training strategies, due to the imbalance of the trojan to non-trojan training data ratio: no weighting, static weighting, and dynamic weighting. Dynamic weighting performed best. Other works address the imbalance in other means, such as synthetic method over-sampling techniques [7]. These optimized methods still do not provide a better understanding of the ML results.
@@ 2.2 Trust-Hub Benchmark Dataset
The Trust-Hub benchmark dataset [8, 9, 10, 6] provides gate-level netlists with known trojan insertions for various circuit types. Some of the Trust-Hub circuits provide gate-level netlists and indicate which gates are part of the trojan. Using these circuits, the trojan gates can be identified from benign gates, those that are legitimate and not part of the trojan. This work uses thirty circuits comprising RS232 communication peripherals and International Symposium on Circuits and Systems (ISCAS) combinational benchmarks. These do not include analog circuit domain or transistor level domains.
@@ 3.1 Machine Learning for Hardware Trojan Detection
Precision computed from their published confusion matrices is 5.13%, indicating that approximately 95% of gates flagged as trojans are false positives. Their published confusion matrices (Table V, dynamic weighting) show an average true negative rate of 49.5%, corresponding to a false positive rate of approximately 50.5%. The sheer volume of false positives limits practical utility for real-world deployment where security engineers must manually inspect falsely flagged gates. This work addresses this limitation by achieving 72.06% precision (95% CI: [61.11%, 82.54%]) at the optimized threshold of 0.980, compared to 11.85% (95% CI: [9.21%, 14.56%]) at the baseline threshold of 0.5, a 14-fold improvement that reduces the false positive burden.
@@ 3.1.2 Recent Advances
In contrast, this work performs gate-level (node-level) trojan localization on heavily imbalanced datasets (approximately 158 to one ratio of benign to trojan gates), directly identifying suspicious gates to enable targeted remediation. While HT-Pred achieves higher overall accuracy on the circuit-level task, the gate-level localization provides the specific information needed for practical trojan removal or mitigation.
@@ 3.1.4 Class Imbalance Handling in Hardware Trojan Detection
Hardware trojan detection faces extreme class imbalance: in typical gate-level netlists, benign gates outnumber trojan gates by ratios exceeding 100 to one, often reaching 158 to one as observed in the Trust-Hub benchmark dataset. Standard machine learning classifiers trained on such imbalanced data exhibit bias toward the majority class, achieving high overall accuracy while failing to detect minority-class trojans.
@@
This work employs XGBoost’s scale_pos_weight parameter to implement cost-sensitive learning following King and Zeng’s weighting approach [12]. By setting the weight ratio equal to the class imbalance ratio of $\frac{N_{benign}}{N_{trojan}}$, the classifier assigns higher loss penalties to trojan misclassifications, encouraging the model to learn discriminative patterns despite extreme rarity. This approach achieves 62.03% recall on 143 to one imbalanced test data, demonstrating that proper imbalance handling is essential for practical trojan detection deployment.
@@ 3.2.2 XAI for Hardware Security
The application of explainable AI to hardware trojan detection remains largely unexplored in published literature. A systematic literature search was conducted using query terms: “explainable AI hardware trojan,” “interpretable machine learning hardware security,” “XAI hardware trojan detection,” and “explainability gate-level netlist analysis.” The search covered publications from 2016 to 2024 (post-LIME and SHAP introduction) and yielded no significant peer-reviewed papers systematically evaluating XAI methods for hardware trojan detection.
Most prior work on ML-based trojan detection uses opaque-box models without explainability analysis [4, 5, 11, 7]. When explanations are provided, they typically consist of simple feature importance rankings extracted from tree-based models (e.g., Random Forest feature importances) without thorough evaluation of explanation quality, consistency across methods, or actionability for security practitioners. Insufficient prior work has compared domain-aware versus model-agnostic explainability techniques, assessed explanation consistency through inter-method correlation, or quantified computational costs of explanation generation at the gate-level. Yet hardware security engineers must validate detection results through manual inspection and design remediation strategies based on identified vulnerabilities. Generic feature importance scores like “LGFi has weight 0.73” provide limited actionability compared to domain-grounded explanations like “high fanin complexity (12 inputs at 2 levels) near primary output matches rare-event trigger signature.”